Mary purchased 100 shares of Sweet Pea Co. stock at a price of $41.51 six months ago. She sold all stocks today for $43.97. During that period the stock paid dividends of $1.38 per share. What is Mary’s effective annual rate?

Round the answers to two decimal places in percentage form.
Answer #1

return of stock = (final price - initial price + dividend)/initial price

= (43.97 - 41.51 + 1.38)/41.51

= 9.25%

Effective annual rate = (1+ APR/n)^n - 1

= (1+ 9.25%)^2 - 1

= 19.36%